OpenStudy (anonymous):

(b^-2/3)^9/2 I'm stumped on this problem. I'm simplifying by the laws of exponents and I can't have any negative exponents. I've got to this point and I'm not sure what to do next (b^-2/3)^9/2=(1/b^2/3)^9/2 any advice?

jhonyy9 (jhonyy9):

you get b^(-3) = 1/b^3

jhonyy9 (jhonyy9):

first is b^(-18/6)

jhonyy9 (jhonyy9):

because (x^a)^b = x^(a*b)

OpenStudy (anonymous):

I think I may typed it in wrong...(b-2/3)9/2 (the fractions are exponents)

jhonyy9 (jhonyy9):

yes this is ok

jhonyy9 (jhonyy9):

so than you get -18/6 like exponent what equal -3

jhonyy9 (jhonyy9):

so b^-3 = 1/b^3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

got it! Thank you so much!!

jhonyy9 (jhonyy9):

yw good luck bye